[Koji, Jenne]
Steve pointed this out to me today, and Koji and I just took a look at it together: The total power coming out of the MOPA box is constant, about 2.7W. However, the NPRO power (as measured by 126MOPA_126MON) has decreased from where we left it last night. It's an exponential decay, and Koji and I aren't sure what is causing it. This may be some misalignment on the PD which actually measures 126MON or something though, because 126MOPA_LMON, which measures the NPRO power inside the NPRO box (that's how it looks on the MEDM screen at least...) has stayed constant. I'm hesitant to be sure that it's a misalignment issue since the decay is gradual, rather than a jump.
Koji and I are going to keep an eye on the 126MON value. Perhaps on Monday we'll take a look at maybe aligning the beam onto this PD, and look at the impedance of both this PD, and the AMPMON PD to see why the reading on the DMM changed last night when we had the DAQ cable T-ed in, and not T-ed in. |